Please help me with number 6,7 and 8Use the information of Mbombela Traders for the year ended 31 March 2020 provided below, to answer questions 6 to 8: Vehicles
Answer #1

Answer 6 :- Option 2 is correct 1,100,000

Explaination

Non current Assets = Vehicles + Equipment

Non current Assets= 700,000 + 400,000

Non current Assets = 1,100,000

Answer 7 :- Option 2 is correct 550,000

Explaination

Current Liability = Bank Overdraft + Creditors control

Current Liability = 50,000 + 500,000

Non current Assets = 550,000

Answer 8 :- Option 4 is correct 1,900,000

Explaination

Total Assets = Vehicles + Inventory + Debtors control + Equipment

Total Assets = 700,000 + 200,000 + 600,000 + 400,000

Total Assets= 1,900,000